OpenStudy (anonymous):

divide the sum of (y+1)^2and (y+1)(y-1) by 2y

OpenStudy (anonymous):

\[\frac{ \left( y+1 \right)^2+\left( y+1 \right)\left( y-1 \right) }{ 2y }=\frac{ \left( y+1 \right)\left( y+1+y-1 \right) }{ 2y }=?\]